### v1

@@ Abstract
We design a suite of minimal algorithmic tasks that are a loose abstraction of open-ended real-world tasks. This allows us to cleanly and controllably quantify the creative limits of the present-day language model. Much like real-world tasks that require a creative, far-sighted leap of thought, our tasks require an implicit, open-ended stochastic planning step that either (a) discovers new connections in an abstract knowledge graph (like in wordplay, drawing analogies, or research) or (b) constructs new patterns (like in designing math problems or new proteins). In these tasks, we empirically and conceptually argue how next-token learning is myopic and memorizes excessively; comparatively, multi-token approaches, namely teacherless training and diffusion models, excel in producing diverse and original output. Secondly, in our tasks, we find that to elicit randomness from the Transformer without hurting coherence, it is better to inject noise right at the input layer (via a method we dub hash-conditioning) rather than defer to temperature sampling from the output layer. Thus, our work offers a principled, minimal test-bed for analyzing open-ended creative skills, and offers new arguments for going beyond next-token learning and softmax-based sampling. We make part of the code available under https://github.com/chenwu98/algorithmic-creativity
@@ 1 Introduction
Not all forms of intelligence are solely about being correct or wrong. In open-ended tasks, what also matters is finding creative ways to satisfy a request, making surprising and fresh connections never seen before. For instance, consider responding to highly under-specified prompts like “Generate a challenging high-school word problem involving the Pythagoras Theorem.” or “Suggest some candidate therapeutic antibodies targeting the HER2 antigen.” or “Provide a vivid analogy to differentiate quantum and classical mechanics.” Creativity in such tasks requires generating responses that are not just correct or coherent, but are also diverse across responses and are original compared to the training data. These currently-sidelined desiderata will rise to prominence as we explore LLMs for open-ended scientific discovery (Gruver et al., 2023; Romera-Paredes et al., 2024; Si et al., 2024; Lu et al., 2024a), for generating novel training data (Yu et al., 2024; Yang et al., 2024c; Wang et al., 2023), and as we scale up test-time compute approaches that benefit from diversity in exploration, such as best-of-N Cobbe et al. (2021); Chow et al. (2024); Dang et al. (2025) and long chain-of-thought reasoning OpenAI (2024); DeepSeek-AI (2025); Snell et al. (2024); Wu et al. (2024).
@@
Concretely, for the scope of this paper, a creative leap-of-thought task refers to tasks that involve a search-and-plan process; crucially, this process orchestrates multiple random decisions in advance before generating the output. Typically, such a leap of thought is highly implicit in the text — to infer it, one has to deeply engage with the text and detect higher-order patterns in it. We think of tasks like designing a satisfying math problem, generating worthwhile research ideas, or drawing surprising analogies as examples of such tasks.
@@
Next, we turn to the de facto approach for randomization in a Transformer: temperature sampling from the output softmax layer. We contrast this against an input-layer randomization approach we call hash-conditioning where we train models with random hash prefixes. We find that, in our tasks, not only does hash-conditioning induce non-trivial algorithmic creativity (even with deterministic, greedy decoding!), hash-conditioning is also competitive with or better than the conventional output-randomization (i.e., temperature sampling). Intuitively, maximizing diversity at the output-token-level is computationally burdensome: it requires simultaneously processing a diverse set of leaps of thoughts to compute a marginalized token distribution. It is easier to first sample a single latent leap of thought, and then compute the token conditioned on that one leap. We conjecture that hash-conditioning enables this conditioned token generation.
@@
As a second direction of progress, we hope our work provides a foundation to think about open-ended tasks which are extremely hard to quantify in the wild. This may spur more algorithmic explorations on improving diversity (such as our approach of hash-conditioning) and on curbing verbatim memorization in language models.
@@
    We find that hash-conditioning i.e., training with random hash prefixes, greatly improves diversity of algorithmic creativity in our tasks, compared to the standard paradigm of temperature sampling.
@@ 2 Open-ended algorithmic tasks & two types of creativity
We are interested in designing simple algorithmic tasks that are loosely inspired by endeavors such as generating scientific ideas, wordplay, narration, or problem-set design, where one needs to generate strings that are both “interesting” and never seen before. In all these tasks, before generating the output, one requires a (creative) leap of thought, a process that (a) is implicit i.e., is not spelled out in token space (or is even inherently hard to spell out), (b) involves discrete random choices (c) and together, those choices must be coherent in that they are carefully planned to satisfy various non-trivial, discrete constraints. These constraints fundamentally define the task and make it interesting e.g., a word problem should be solvable by arithmetic rules, or a pun must deliver a surprising punchline. The goal in such open-ended tasks is not just coherence though, but also diversity and novelty — generations must be as varied as possible and must not be regurgitated training data. Before we design tasks that capture the aforementioned leap of thought, we first clarify what tasks do not require such a step.
@@ 2.1 The fundamental types of creativity (Boden, 2003)
Combinational creativity. [^1] Some call it combinatorial creativity. We use the term from Boden (2003), combinational. Consider rudimentary wordplay of the form “What musical genre do balloons enjoy? Pop music.” or “What kind of shoes do spies wear? Sneakers.” There is a global structure here: two unrelated entities (genre & balloons) are related eventually through a punchline (pop); the punchline itself is a mutual neighbor on a semantic graph. More broadly, Boden (2003) argues that many tasks, like the above, involve “making unfamiliar combinations of familiar ideas” or the “unexpected juxtaposition of [known] ideas”. Other tasks include drawing analogies, or finding connections between disparate ideas in science. [^2] Even this very paper’s idea draws a connection between the existing ideas of multi-token prediction, limits of next-token prediction and creative planning tasks. All these tasks involve a leap of thought that in effect searches and plans over a space of known facts and combines them.
@@
In the upcoming sections, we will attempt to capture some (not all) core computational aspects of rudimentary instances within the two classes of creative skills above. We emphasize that by no means does our minimal algorithmic setup intend to capture the human values that go into these endeavors; nor do they capture the rich array of creative acts that Boden (2003) discusses within these categories. (See limitations in §6)
@@ 2.2 The basic setting and notations
Our setup models an in-distribution form of novelty as it offers a rigorous and tractable way to study the problem. Admittedly though, this is a far simpler form of novelty than what is expected in real-world tasks. Nevertheless, even this simple setting will help us appreciate the limits of current language models.
@@ 2.3 Tasks inspired by combinational creativity
Combinational creativity requires a recall-and-search through entities from memory subject to the constraint that they relate to each other in an interesting way. We abstract this through tasks that discover structures from an in-weights graph i.e., a graph stored in the model weights, not reveal in context.
@@ 2.3.1 Sibling discovery
This task involves an implicit, bipartite $\mathcal{G}$ made of parent vertices $\mathscr{V}=\{A,B,C,\ldots\}$ each neighboring a corresponding set of children $\texttt{nbr}(A)=\{a_{1},a_{2},\ldots,\}$, $\texttt{nbr}(B)=\{b_{1},b_{2},\ldots\}$ and so on. We define $\texttt{coh}(\bm{s})$ to hold true on sibling-parent triplets of the form $\bm{s}=(\gamma,\gamma^{\prime},\Gamma)$ such that $\gamma,\gamma^{\prime}\in\texttt{nbr}(\Gamma)$. We then consider a uniform distribution $\mathcal{D}$ over all coherent strings $(\gamma,\gamma^{\prime},\Gamma)$ for a fixed graph $\mathcal{G}$. The model witnesses i.i.d samples from this distribution. During test-time, the model must maximize algorithmic creativity (Eq 1) by generating novel parent-sibling triplets based on its in-weights knowledge of $\mathcal{G}$. Note that the model is not provided the graph in-context as this would sidestep a core computational step in combinational creativity: recalling facts from a large memory (see §B.2). The hope is that the model infers and stores the pairwise adjacencies of $\mathcal{G}$ in its weights (given sufficient data). Full dataset description is in §C and Fig 9.
@@ 2.3.2 Triangle discovery
Next, we design a task that requires a more complex, higher-order planning: generating triangles from an appropriately-constructed knowledge graph $\mathcal{G}=({V},E)$ (which contains many triangles; see §C). Thus, in this task $\texttt{coh}((v_{1},v_{2},v_{3}))=\texttt{true}$ iff all three edges between $\{v_{1},v_{2},v_{3}\}$ belong in $\mathcal{G}$. Furthermore, we define $\texttt{uniq}(\cdot)$ and $\texttt{mem}(\cdot)$ such that various permutations of the same triangle are counted as one (see details in §C, including the exact formatting of the string). Note that the leap of thought in this task is much harder to learn and execute as it requires co-ordinating three edges in parallel, from memory.
@@ 2.4 Tasks inspired by exploratory creativity
Recall that we are also interested in creativity that involves constructing new structures. For instance, this may be designing word problems that correspond to novel solutions. Below, we capture this through tasks that construct adjacency lists of structured graphs. Note that no knowledge graph is involved in these tasks.
@@ 2.4.1 Circle construction
In this task, the generated strings must be randomized adjacency lists that can be rearranged to recover circle graphs of $N$ vertices. Let the generated list be $\bm{s}=(v_{i_{1}},v_{i_{2}}),(v_{i_{3}},v_{i_{4}}),\ldots$. We define $\texttt{coh}(\bm{s})=\texttt{true}$ iff there exists a resolving permutation $\pi$ such that $\pi(\bm{s})=(v_{j_{1}},v_{j_{2}}),(v_{j_{2}},v_{j_{3}}),\ldots(v_{j_{n}},v_{j_{1}})$ for distinct $j_{1},j_{2},\ldots j_{n}$. i.e., each edge leads to the next, and eventually circles back to the first vertex. We define uniq and mem such that different examples with the same resolving $\pi$ are counted as the same, even if they have differing vertices. As always, the learner is then exposed to a finite set of uniformly sampled coherent strings. Note that the latent leap of thought here requires constructing a novel permutation $\pi$ before generating the sequence.
@@ 2.4.2 Line construction
A simple variant of the above task is one where the edge set corresponds to a line graph. The resolving permutation $\pi$ is such that $\pi(\bm{s})=(v_{j_{1}},v_{j_{2}}),(v_{j_{2}},v_{j_{3}})\ldots,(v_{j_{n-1}},v_{j_{n}})$ for distinct $j_{1},j_{2},\ldots j_{n}$. i.e., each edge leads to the next until a dead-end.
@@ 2.5 Permutation-invariance of our tasks
We emphasize a key novel aspect in our last three tasks. Many algorithmic tasks in literature like addition (Lee et al., 2024), or the path-star task (B&N’24) or Sibling Discovery have a natural ordering in which the tokens can be learned and generated, even if it may not be left to right. However, the Triangle Discovery, Line Construction and Circle Construction tasks are permutation-invariant — no token is more privileged than the other, and hence all tokens must be “simultaneously learned” to infer the underlying process. Intuitively, we view this as an abstraction of real-world tasks where the creative process is highly implicit, and not immediate from the text. These tasks offer a test-bed even for non-next-token approaches that rely on re-permuting the tokens (Pannatier et al., 2024; Thankaraj et al., 2025) or predicting only a part of the future (Kitouni et al., 2024; Nolte et al., 2024; Bavarian et al., 2022; Fried et al., 2022).
@@ 3 Training and Inference
For our next-token-trained (NTP) models, we use the standard teacher-forcing objective used in supervised finetuning. Given prompt $\bm{p}$ and ground truth sequence $\bm{s}$, the model is trained to predict the $i$’th token $s_{i}$, given as input the prompt and all ground truth tokens up until that point, $(\bm{p},\bm{s}_{<i})$. We write the objective more explicitly in §A Eq 2.
For the multi-token Transformer models, we use teacherless training (Monea et al., 2023; Bachmann & Nagarajan, 2024; Tschannen et al., 2023), where the model is trained to predict $s_{i}$ simultaneously for all $i$, only given the prompt $\bm{p}$ (and some dummy tokens in place of the $\bm{s}$ that was once given as input). Since the exact details of this is irrelevant to our discussion, we describe this in Eq 2. To train our models, we use a hybrid of this objective and the next-token objective.
@@
3.1 Hash-conditioning for Transformers
Unlike closed-ended tasks where a prompt (i.e., a prefix) maps to a unique correct answer, our open-ended tasks are prompt-free. For a prompt-free autoregressive Transformer to provide diverse outputs, we must use temperature sampling rather than greedy decoding. However, as we show later, training a prompt-free Transformer model even on our simple tasks leads to poor creativity in some of our settings (while this was no problem for our Diffusion models!). As a natural alternative to this, we tried was to prepend a prompt of pause tokens (Goyal et al., 2024) to all datapoints — both during training and during inference — in order to allow extra computation to the model before it emits its outputs.
Next, we tried an even more sophisticated alternative we call as hash-conditioning. Here, we use as prompt, a random hash string unique to each training datapoint (rather than the same constant sequence of pause tokens); during test-time, we prompt with novel hash strings to extract the test data. We provide possible intuitions for why this may help in §5.1.
@@ 4 Experimental results
Part of our experiments are performed for a Gemma v1 (2B) pre-trained model (Gemma Team et al., 2024), averaged over 4 runs. For diffusion, we use a 90M (non-embedding) parameters Score Entropy Discrete Diffusion model (SEDD; Lou et al., 2023). For a fair comparison against NTP, we use a 86M (non-embedding) parameters GPT-2 model Radford et al. (2019).
In all our experiments, we finetune the models until it is clear that algorithmic creativity (Eq. 1) has saturated. All values are reported from this checkpoint. Finally, since our best Transformer results were under hash-conditioning (for both next- and mult-token training), our main results are reported under that training setting; we provide various ablations without that as well. Please see §D for more experimental details, and §C for precise dataset details (e.g., how the graph is constructed, how sequences are formatted etc.,).

### v2

@@ Abstract
We design a suite of minimal algorithmic tasks that are a loose abstraction of open-ended real-world tasks. This allows us to cleanly and controllably quantify the creative limits of the present-day language model. Much like real-world tasks that require a creative, far-sighted leap of thought, our tasks require an implicit, open-ended stochastic planning step that either (a) discovers new connections in an abstract knowledge graph (like in wordplay, drawing analogies, or research) or (b) constructs new patterns (like in designing math problems or new proteins). In these tasks, we empirically and conceptually argue how next-token learning is myopic and memorizes excessively; multi-token approaches, namely teacherless training and diffusion models, comparatively excel in producing diverse and original output. Secondly, to elicit randomness without hurting coherence, we find that injecting noise at the input layer (dubbed as seed-conditioning) works surprisingly as well as (and in some conditions, better than) temperature sampling from the output layer. Thus, our work offers a principled, minimal test-bed for analyzing open-ended creative skills, and offers new arguments for going beyond next-token learning and temperature sampling. We make part of the code available under https://github.com/chenwu98/algorithmic-creativity
next-token prediction, multi-token prediction, creativity
@@ 1 Introduction
Not all forms of intelligence are solely about being correct or wrong. In open-ended tasks, what also matters is finding creative ways to satisfy a request, making surprising and fresh connections never seen before. For instance, consider responding to highly under-specified prompts like “Generate a challenging high-school word problem involving the Pythagoras Theorem.” or “Suggest some candidate therapeutic antibodies targeting the HER2 antigen.” or “Provide a vivid analogy to differentiate quantum and classical mechanics.” Creativity in such tasks requires generating responses that are not just correct or coherent, but are also diverse across responses and are original compared to the training data. These currently-sidelined desiderata will rise to prominence as we explore LLMs for open-ended scientific discovery (Gruver et al., 2023; Romera-Paredes et al., 2024; Si et al., 2024; Lu et al., 2024a), for generating novel training data (Yu et al., 2024; Yang et al., 2024c; Wang et al., 2023), and as we scale up test-time compute approaches that benefit from diversity in exploration, such as best-of-N (Cobbe et al., 2021; Chow et al., 2024; Dang et al., 2025) and long chain-of-thought reasoning (OpenAI, 2024; DeepSeek-AI, 2025; Snell et al., 2024; Wu et al., 2024).
@@
Concretely, for the scope of this paper, a creative leap-of-thought task refers to tasks that involve a search-and-plan process; crucially, this process orchestrates multiple random decisions in advance before generating the output. Typically, such a leap of thought is highly implicit in the text — to infer it, one has to deeply engage with the text and detect higher-order patterns. We can think of tasks like designing satisfying math problems, generating worthwhile research ideas, or drawing surprising analogies as examples of such tasks.
@@
Next, we turn to how we elicit randomness from a Transformer. While the de facto approach is to elicit randomness at the output — temperature sampling — we contrast this with injecting randomness into the model. Concretely, we study seed-conditioning, where we train with a random prefix string per example. Surprisingly, we find that seed-conditioning induces non-trivial algorithmic creativity even with deterministic, greedy decoding. In fact, seed-conditioning is competitive with and in some cases, more creative than temperature sampling. Intuitively, maximizing diversity at the output-token-level is computationally burdensome: it requires simultaneously processing a diverse set of leaps of thoughts to compute a marginalized token distribution. It is easier to first sample a single latent leap of thought, and then compute the token conditioned on that one leap. We conjecture that seed-conditioning enables a latent-leap-conditioned token generation.
@@
As a second direction of progress, we hope our work provides a foundation to think about open-ended tasks which are extremely hard to quantify in the wild. This may spur more algorithmic explorations on improving diversity (such as our approach of seed-conditioning) and on curbing verbatim memorization in language models.
@@
    We find that input-randomization via seed-conditioning achieves algorithmic creativity competitive with conventional output-randomization via temperature sampling.
@@ 2 Open-ended algorithmic tasks & two types of creativity
We are interested in designing simple algorithmic tasks that are loosely inspired by endeavors such as generating scientific ideas, wordplay, narration, or problem-set design, where one needs to generate strings that are both “interesting” and never seen before. In all these tasks, before generating the output, one requires a (creative) leap of thought, a process that (a) is implicit i.e., is not spelled out in token space (or is even inherently hard to spell out), (b) involves discrete random choices (c) and together, those choices must be coherent in that they are carefully planned to satisfy various non-trivial, discrete constraints. These constraints fundamentally define the task and make it interesting e.g., a word problem should be solvable by arithmetic rules, or a pun must deliver a surprising punchline. The goal in such open-ended tasks is not just coherence though, but also diversity and novelty — generations must be as varied as possible and must not be regurgitated training data. Before we design tasks that require the aforementioned leap of thought, we first clarify what tasks do not require it.
@@ 2.1 The fundamental types of creativity (Boden, 2003)
Combinational creativity. Consider rudimentary wordplay of the form “What musical genre do balloons enjoy? Pop music.” or “What kind of shoes do spies wear? Sneakers.” There is a global structure here: two unrelated entities (genre & balloons) are related eventually through a punchline (pop); the punchline itself is a mutual neighbor on a semantic graph. More broadly, Boden (2003) argues that many tasks, like the above, involve “making unfamiliar combinations of familiar ideas” or the “unexpected juxtaposition of [known] ideas”. Other tasks include drawing analogies, or finding connections between disparate ideas in science. All these tasks involve a leap of thought that in effect searches and plans over a space of known facts and combines them.
@@
In the upcoming sections, we will attempt to capture some core computations of rudimentary instances within the two classes of creative skills above. We emphasize that by no means does our minimal algorithmic setup intend to capture the human values that go into these endeavors; nor do they capture the rich array of creative acts that Boden (2003) discusses within these categories. (See limitations in §6)
@@ 2.2 The basic setting and notations
Our setup models an in-distribution form of novelty as it offers a rigorous way to study the problem. Admittedly, this is a far simpler form of novelty than what is expected in the real-world. Nevertheless, even this simple setting will help us appreciate the limits of current language models.
@@ 2.3 Tasks inspired by combinational creativity
Combinational creativity requires a recall-and-search through entities from memory, constrained to relate with each other in an interesting way. We abstract this through tasks that discover structures from an in-weights graph i.e., a graph stored in the model weights, not reveal in context.
@@ 2.3.1 Sibling discovery
This task involves an implicit, bipartite $\mathcal{G}$ with parent vertices $\mathscr{V}=\{A,B,C,\ldots\}$ each neighboring a set of children $\texttt{nbr}(A)=\{a_{1},a_{2},\ldots,\}$, $\texttt{nbr}(B)=\{b_{1},b_{2},\ldots\}$ and so on. We define $\texttt{coh}(\bm{s})$ to be true on sibling-parent triplets of the form $\bm{s}=(\gamma,\gamma^{\prime},\Gamma)$ such that $\gamma,\gamma^{\prime}\in\texttt{nbr}(\Gamma)$. We then consider a uniform distribution $\mathcal{D}$ over all coherent strings $(\gamma,\gamma^{\prime},\Gamma)$ for a fixed graph $\mathcal{G}$. The model witnesses i.i.d samples from $\mathcal{D}$. During test-time, the model must maximize algorithmic creativity (Eq 1) by generating novel parent-sibling triplets based on its in-weights knowledge of $\mathcal{G}$. Note that the model is not provided the graph in-context as this would sidestep a core computational step in combinational creativity: recalling facts from a large memory (see §B.2). The hope is that the model infers and stores the pairwise adjacencies of $\mathcal{G}$ in its weights (given sufficient data). Full dataset description is in §C and Fig 9.
@@ 2.3.2 Triangle discovery
Next, we design a task that requires a more complex, higher-order planning: generating triangles from an appropriately-constructed knowledge graph $\mathcal{G}=({V},E)$ (see graph construction in §C). Thus, in this task $\texttt{coh}((v_{1},v_{2},v_{3}))=\texttt{true}$ iff all three edges between $\{v_{1},v_{2},v_{3}\}$ belong in $\mathcal{G}$. Furthermore, we define $\texttt{uniq}(\cdot)$ and $\texttt{mem}(\cdot)$ such that various permutations of the same triangle are counted as one (see details in §C, including the exact formatting of the string). Note that the leap of thought in this task is much harder to learn and execute as it requires co-ordinating three edges in parallel, from memory.
@@ 2.4 Tasks inspired by exploratory creativity
We are also interested in the less-knowledge-intensive creativity involved in tasks like designing word problems that demand novel solutions. We capture this skill minimally through tasks that construct new structures. No knowledge graph is involved in these tasks.
@@ 2.4.1 Circle construction
In this task, the model must generate adjacency lists that can be rearranged to recover circle graphs of $N$ vertices. Let the generated list be $\bm{s}=(v_{i_{1}},v_{i_{2}}),(v_{i_{3}},v_{i_{4}}),\ldots$. We define $\texttt{coh}(\bm{s})=\texttt{true}$ iff there exists a resolving permutation $\pi$ such that $\pi(\bm{s})=(v_{j_{1}},v_{j_{2}}),(v_{j_{2}},v_{j_{3}}),\ldots(v_{j_{n}},v_{j_{1}})$ for distinct $j_{1},j_{2},\ldots j_{n}$. i.e., each edge leads to the next, and eventually circles back to the first vertex. We define uniq and mem such that different examples with the same resolving $\pi$ are counted as the same, even if they have differing vertices. As always, the learner is then exposed to a finite set of uniformly sampled coherent strings. Note that the latent leap of thought here requires constructing a novel permutation $\pi$ before generating the sequence.
@@ 2.4.2 Line construction
A simple variant of the above task is one where the edge list is of a line graph. The resolving permutation $\pi$ satisfies $\pi(\bm{s})=(v_{j_{1}},v_{j_{2}}),(v_{j_{2}},v_{j_{3}})\ldots,(v_{j_{n-1}},v_{j_{n}})$ for distinct $j_{1}\ldots j_{n}$. i.e., each edge leads to the next until a dead-end.
@@ 2.5 Permutation-invariance of our tasks
We emphasize a key novel aspect in our last three tasks. Many algorithmic tasks in literature like addition (Lee et al., 2024), or the path-star task (B&N’24) or Sibling Discovery have a natural ordering in which the tokens can be learned and generated, even if not left to right. However, the Triangle Discovery, Line Construction and Circle Construction tasks are permutation-invariant — no token is more privileged than the other, and all tokens must be “simultaneously learned” to infer the underlying process. We view this as an abstraction of real-world tasks where the creative process is highly implicit, and not immediate from the text. These tasks offer a test-bed even for non-next-token approaches that rely on re-permuting the tokens (Pannatier et al., 2024; Thankaraj et al., 2025) or predicting only a part of the future (Kitouni et al., 2024; Nolte et al., 2024; Bavarian et al., 2022; Fried et al., 2022).
@@ 3 Training and Inference
For our next-token-trained (NTP) models, we use the standard teacher-forcing objective used in supervised finetuning. Given prompt $\bm{p}$ and ground truth sequence $\bm{s}$, the model is trained to predict the $i$’th token $s_{i}$, given as input the prompt and all ground truth tokens up until that point, $(\bm{p},\bm{s}_{<i})$. We write the objective more explicitly in §A Eq 2. For the multi-token Transformer models, we use teacherless training (Monea et al., 2023; Bachmann & Nagarajan, 2024; Tschannen et al., 2023), where the model is trained to predict $s_{i}$ simultaneously for all $i$, only given the prompt $\bm{p}$ (and some dummy tokens in place of the $\bm{s}$ that was once given as input). Since the exact details of this is irrelevant to our discussion, we describe this in Eq 2. To train our models, we use a hybrid of this objective and the next-token objective.
@@
3.1 Seed-conditioning
Unlike closed-ended tasks where a prompt (i.e., a prefix) maps to a unique correct answer, our open-ended tasks are prompt-free. As an alternative to this, we also tried prepending a prompt of pause tokens (Goyal et al., 2024) to all datapoints — both during training and during inference — to enable extra computation before the model emits its outputs. Next, we tried a more sophisticated alternative we term seed-conditioning. Here, we associate each training point with an arbitrary (meaningless) prefix to the model (e.g., of random characters) called a seed; during test-time, we prompt with novel seeds hoping to extract novel test data. We provide hypotheses motivating this procedure in §5.1.
@@ 4 Experimental results
Part of our experiments are performed for a Gemma v1 (2B) pre-trained model (Gemma Team et al., 2024), averaged over 4 runs. For diffusion, we use a 90M (non-embedding) parameters Score Entropy Discrete Diffusion model (SEDD; Lou et al., 2023). For a fair comparison against NTP, we use a 86M (non-embedding) parameters GPT-2 model (Radford et al., 2019). In all our experiments, we finetune the models until it is clear that algorithmic creativity (Eq. 1) has saturated. All values are reported from this checkpoint. Finally, since our best Transformer results were under noise-conditioning (for both next- and mult-token training), our main results are reported under that training setting; we provide various ablations without that as well. Please see §D for more experimental details, and §C for precise dataset details (e.g., how the graph is constructed, how sequences are formatted etc.,).